# Why We Need More than General Multimodal LLMs

### v1

@@ Appendix A Quantitative Comparison on Music AVQA Datasets
Att-BLSTM [132] 60.00 49.55 54.77 32.15 47.97 48.89 54.33 39.46 32.52 51.00 24.45 40.35 40.35
AVSD [96] 50.92 54.20 52.56 35.21 68.11 52.20 64.13 36.68 27.14 58.99 40.83 45.55 45.55
CONVLSTM [32] 55.68 60.22 57.95 35.64 51.66 52.23 72.45 53.18 32.35 57.91 43.33 51.84 51.84
FCNLSTM [32] 51.36 57.96 54.66 33.53 52.96 50.09 71.64 51.98 34.96 57.40 33.90 49.98 49.98
GRU [4] 57.78 58.95 58.36 38.08 57.67 54.17 70.53 43.33 39.70 57.29 35.85 49.34 49.34
HCAttn [82] 51.65 53.12 52.38 32.86 60.09 50.02 63.85 39.77 36.01 54.47 36.54 46.13 46.13
HCRN [65] 54.42 39.81 47.11 32.71 45.34 43.88 53.63 39.67 37.08 35.10 42.30 41.56 41.56
HME [31] 58.28 56.63 57.45 33.71 65.93 54.40 66.12 39.91 40.18 56.89 37.76 48.17 48.17
LAVisH [76] 52.86 62.72 57.79 38.33 67.47 55.83 78.65 41.48 32.38 62.18 44.05 51.75 51.75
LAViT [123] 47.01 47.86 47.43 31.39 66.35 48.01 37.21 53.02 36.87 43.05 42.17 42.46 42.46
MCAN [121] 67.59 54.49 61.04 45.64 64.37 58.62 59.29 53.86 45.02 51.49 46.35 51.20 51.20
MCCD [86] 75.78 63.43 69.60 61.76 73.43 68.80 76.18 50.55 50.92 62.15 66.95 61.35 61.35
PSAC [71] 54.85 52.77 53.81 37.99 66.83 53.25 53.05 47.14 38.14 48.53 36.46 44.66 44.66
@@ Appendix C Representative Examples of Music AVQA Question Types
Appendix D How Music AVQA Differs from Traditional Multimodal Tasks
Recent progress in multimodal research has led to a wide range of models, including vision–language pretraining [127, 126], video captioning [56], instructional video QA [70] and mathematical reasoning [44], multimodal content generation and dialogue [36, 49, 22, 57], scene-level retrieval [48], audio–visual representation learning [40, 23], multimodal graph-based classification [92], multimodal clustering [118, 117], multi-label classification [112, 113], and symbolic theory-of-mind reasoning [128]. Among these advances, Music AVQA stands out within audio–visual–text (AVT) tasks for requiring models to disentangle polyphonic, continuous sound streams, bind them to precise visual sources, and reason with explicit musical knowledge.
Generic Audio–Visual Question Answering.
Classic AVQA benchmarks [37, 9, 102, 131] present short clips with a single audible event (e.g., a dog bark) and ask coarse “what/where/when” questions that can be answered once the sounding object is localized and its time span identified. The AVQA dataset itself illustrates this design: most videos contain one foreground sound and minimal polyphony, so millisecond-level alignment is unnecessary [69, 116]. Music AVQA, in contrast, poses queries such as “Which violin enters after the flute?”, which requires the model to track multiple overlapping instruments and reason over precise temporal order, a level of granularity generic AVQA never targets.
Video Captioning.
Datasets like MSR-VTT [114], GIF [72], ActivityNet Captions [62], How2 [95], and Vatex [107] evaluate whether a system can produce one or two fluent sentences that summarize the gist of a clip. Moreover, temporal mis-alignment of a few seconds or the omission of background sounds rarely affects the score. Music AVQA eliminates summarisation entirely and replaces it with beat-accurate, question-driven reasoning: the model must pinpoint note onsets, match them to performers, and compare rhythmic or dynamic patterns tasks far beyond the scope of generic captioning.
Instructional Video QA.
TutorialVQA [19], YouCookQA [106] and HowToVQA69M [115] frame question answering around narrated procedural videos whose audio channel is dominated by speech that explicitly describes each step. The narration acts as a guide-track and overlapping non-speech sounds are rare and not queried. Music AVQA removes narration altogether and treats the dense musical audio as the primary reasoning target, forcing models to infer structure (beats, phrases, instrument entrances) directly from raw sound rather than from textual guidance.
Audio–Visual Scene-Aware Dialogue.
The AVSD dataset [2] and its follow-up [3, 96, 28, 97] train systems to hold multi-turn conversations about short ($\approx$ 10 s) household videos, grounding answers in coarse scene context while maintaining dialog coherence. Acoustic events are typically brief (speech, clatter) and alignment at ±$\pm 1$ s suffices. Music AVQA disregards conversational flow and instead demands hierarchical timing: beats, bars, sections. Every answer depends on tight audio–video synchrony, not on dialog history management. Recent studies on diversified and persona-aware dialogue generation [66, 67] further underscore the importance of fine-grained temporal and semantic alignment—even in unimodal dialogue settings.
Multimodal Sentiment Analysis.
Benchmarks such as CMU-MOSI [124] and CMU-MOSEI [104] fuse text transcripts, facial expression and prosody to predict sentiment or emotion over 5 to 30 second clips. Overlap between speakers or sound sources is noise to suppress, and no explicit source-binding is required. In Music AVQA overlap is the signal: models must isolate each instrument’s contribution, bind it to its on-screen performer and reason about their relationships (loudness, order, count). Other emotion datasets, like IEMOCAP [8], MELD [91], and CH-SIMS [120], use 5–30 s clips, treat overlapping voices as background noise, and label sentiment at the utterance—rather than source—level.
Cross-Modal Retrieval.
Contrastive systems such as CLIP [93] (vision–text), ALIGN [55], LiT [125], or ImageBind [38] (audio–vision–text) learn global embeddings and judge success by top-$k$ similarity—minor temporal or spatial errors barely change the score. Music AVQA instead penalises any mis-alignment: swapping left/right instruments or missing a single beat flips the answer. The task therefore demands persistent token-level grounding rather than coarse embedding proximity.
@@ E.1 Music AVQA Datasets
Table 6: Evolution and characteristics of Music AVQA datasets: a comparative overview of MUSIC-AVQA [69], MUSIC-AVQA v2.0 [80], and MUSIC-AVQA-R [86] benchmarks.
@@
MUSIC-AVQA [69] The MUSIC-AVQA dataset represents a significant contribution to audio-visual question answering research, comprising 9,288 videos with over 150 hours of musical performances covering 22 instruments, generating 45,867 question-answer pairs. The dataset is randomly split into training, validation, and testing sets with 32,087, 4,595, and 9,185 QA pairs respectively, spanning 33 question templates across 9 question types including existential, location, counting, comparative, and temporal questions.
MUSIC-AVQA v2.0 [80] The MUSIC-AVQA v2.0 dataset builds upon the original MUSIC-AVQA by addressing data bias issues, comprising 10,518 videos (9,288 from the original plus 1,230 new videos) with musical performances covering 22 instruments, generating approximately 54,000 question-answer pairs. The balanced dataset splits into training, validation, and testing sets with 36,700, 5,250, and 10,819 QA pairs respectively, spanning 33 question templates across 9 question types. The authors specifically balance 15 biased templates by ensuring no dominant answers exceed 60% for binary questions or 50% for multi-class questions, particularly enhancing representation of underrepresented answers in existential, counting, temporal, location, and comparative question categories.
MUSIC-AVQA-R [86] The MUSIC-AVQA-R dataset proposed in this paper is an extension of MUSIC-AVQA specifically designed to evaluate the robustness of audio-visual question answering models. It expands the original test set through a human-machine collaboration mechanism that rephrases each question 25 times, increasing the number of questions from 9,129 to 211,572, and introduces distribution shifts to categorize questions into head (common) and tail (rare) samples. Compared to the original dataset, MUSIC-AVQA-R features a vocabulary size of 465 (five times larger than MUSIC-AVQA), provides more diverse question formulations while preserving inherent biases in the training and validation sets, and offers three evaluation metrics—head accuracy, tail accuracy, and overall accuracy—enabling researchers to assess model performance in both in-distribution and out-of-distribution scenarios, making it the first dataset specifically designed for robustness evaluation in audio-visual question answering tasks.
@@ E.2 Key Difference from Other AVQA Datasets
Table 7 contrasts the Music-AVQA dataset [69] with several widely-used AVQA benchmarks [116, 87, 111, 78, 9]. For each dataset, we highlight the most salient divergence from the music-specific setting, focusing on aspects such as task format, content domain, temporal scope, and the presence or absence of fine-grained musical reasoning.
Table 7: Other representative benchmarks (AVQA [116], EgoSchema [87], FunQA [111], VALOR-1M [78], and VGG-Sound [9]) and the key difference each bears with respect to Music-AVQA [69].
@@
AVQA [116] Builds multiple-choice QA on everyday VGG-Sound clips; questions target generic activities and causal relations in real-life videos, so it lacks the fine-grained instrument/sound localization and music-theory knowledge required in MUSIC-AVQA.
@@
EgoSchema [87] Uses first-person (Ego4D) footage that is three-minutes long, stressing long-range temporal reasoning in egocentric daily tasks; audio cues are incidental and the task is 5-way multiple choice, very different from the short, professionally filmed music performances and open-ended answers in MUSIC-AVQA.
@@
FunQA [111] Focuses on“surprising” humour/creative/magic clips (4.3 k videos, 312 k QAs) that test commonsense violations; audio is often background and questions centre on counter-intuitive visual events, not on synchronised musical notes or instrument semantics.
@@
VALOR-1M [78] A pre-training corpus (1 M videos) with tri-modal captions meant for retrieval/captioning; QA supervision is extremely sparse and relies on auto-generated subtitles, so it serves as a foundation model resource rather than a targeted AVQA evaluation set like MUSIC-AVQA.
@@
VGG-Sound [9] It is an audio-visual correspondent dataset consisting of short clips of audio sounds from YouTuBe. And it provides raw audio–visual correspondence but no question–answer supervision or fine-grained reasoning labels.
@@
Appendix F Details of Music AVQA Methods with Spatial-Temporal Designs
@@
Amuse [24] Focuses on music performance scenarios by aligning time segments in both audio and video streams via a cross-attention paradigm. Exploits synchronized features (such as beat-level or note-level alignment) to capture subtle temporal dependencies among instruments in dense music passages. By integrating rhythmic cues and cross-modal interactions, it is particularly suited for questions that involve multiple instruments playing simultaneously or changing their patterns over time.
@@
AVST [69] Proposes a spatio-temporal grounded audio-visual approach that explicitly localizes sounding objects in each frame while applying a question-guided temporal attention mechanism. The model grounds audio-visual events and emphasizes which frames (visual) and which segments (audio) are most relevant for question answering. By combining localized visual features and temporal cues, it captures object interactions over time and can better handle questions involving spatial and temporal reasoning.
@@
CIGN [88] Learns audio-visual class tokens and an Audio-Visual Continual Grouping module that, at every time-step, pulls together frame-level spectrogram features and region features into category-aware clusters. A token-distillation schedule preserves past knowledge while the regrouping logic tracks objects and sounds through the video’s timeline, giving the model temporally consistent, cross-modal semantics for spatial-temporal reasoning.
@@
DCL [84] Introduces a Disentangled Counterfactual Learning framework to handle physical audio-visual commonsense reasoning tasks. Decomposes video signals into static (time-invariant) and dynamic (time-varying) factors using a VAE-based encoder, enabling clearer separation of constant background features from changing events. Additionally employs a counterfactual intervention module on the dynamic factors to perform causal reasoning, helping the model answer “what if” questions related to temporal and event relationships.
@@
DG-SCT [27] Introduces a Dual-Guided Spatial-Channel-Temporal (DG-SCT) attention layer that is injected in every frozen audio and visual transformer block. Audio prompts steer visual tokens (and vice-versa) via bidirectional attention that highlights salient spatial regions, discriminative feature channels, and pivotal temporal segments, producing fine-grained spatio-temporal alignments that boost related tasks.
@@
EEMC [108] Divides audio/video into 1-s slices and fuses them with text through a Temporal Bi-modal Transformer backed by a cached-memory mechanism that magnifies sudden changes across time. The resulting multimodal cue stream then serves as a cross-attention prompt for the segmentation decoder, enabling precise localisation of objects and events as their spatial footprints and temporal order evolve.
@@
LAST-Att [80] Tackles audio-visual question answering with a repeated cross-attention strategy. Uses Swin-Transformer-v2 for visual frame features and a specialized Audio Spectrogram Transformer for audio, then merges them based on the question. By repeatedly “attending” to the most relevant frames and spectrogram patches, it effectively localizes musical actions (e.g., a pianist’s keystrokes) over time. This design is well suited for intricate temporal queries and locating key audio events in dense musical content.
@@
LAVisH [76] Adds a lightweight Latent Audio-Visual HYbrid adapter to every layer of a frozen ViT. A compact pool of latent tokens acts as a cross-attention bottleneck, letting audio frames gate visual tokens (and vice-versa) as the video unfolds, so spatial patches and framewise dynamics are fused early while keeping the backbone frozen.
@@
LAViT [123] Targets 360° videos with a transformer that augments each patch by a quaternion-based spherical coordinate and aligns it with audio via joint contrastive objectives. The spherical embedding plus an auxiliary audio-skewness prediction head lets the model reason about where (on the sphere) and when a sound arises, delivering fine-grained spatial-temporal grounding beyond normal FOV clips.
@@
LSTTA [77] A parameter-efficient transfer learning approach for audio-visual-language tasks that adds dedicated adapter modules while freezing large pretrained backbones. Splits temporal modeling into two scales: a short-term semantic interaction module (for capturing local correlations such as brief instrumental flourishes) and a long-term semantic filtering module (for broader progressions over many frames). This structure helps the model identify when, how, and for how long different instruments contribute, achieving a refined spatio-temporal representation.
@@
MAVEN [85] Employs a Multimodal Audio-Visual Epistemic Network that cycles between audio, video and text logits, using debiasing constraints to keep modality-specific and fused predictions consistent over time. The cycle guidance implicitly anchors each question to the correct temporal segments while suppressing spurious correlations.
@@
MCCD [86] Introduces a Multifaceted Cycle-Collaborative Debiasing objective: KL penalties enlarge the gap between uni-modal and tri-modal logits at every timestep, then force the three unimodal paths to agree with each other. This temporal-cycle training steers attention toward frames (and sounds) that all modalities truly share, yielding stabler spatial-temporal grounding under distribution shift.

### v2

@@ Appendix A Quantitative Comparison on Music AVQA Datasets
Att-BLSTM Zhou et al. (2016) 60.00 49.55 54.77 32.15 47.97 48.89 54.33 39.46 32.52 51.00 24.45 40.35 40.35
AVSD Schwartz et al. (2019) 50.92 54.20 52.56 35.21 68.11 52.20 64.13 36.68 27.14 58.99 40.83 45.55 45.55
CONVLSTM Fayek and Johnson (2020) 55.68 60.22 57.95 35.64 51.66 52.23 72.45 53.18 32.35 57.91 43.33 51.84 51.84
FCNLSTM Fayek and Johnson (2020) 51.36 57.96 54.66 33.53 52.96 50.09 71.64 51.98 34.96 57.40 33.90 49.98 49.98
GRU Antol et al. (2015) 57.78 58.95 58.36 38.08 57.67 54.17 70.53 43.33 39.70 57.29 35.85 49.34 49.34
HCAttn Lu et al. (2016) 51.65 53.12 52.38 32.86 60.09 50.02 63.85 39.77 36.01 54.47 36.54 46.13 46.13
HCRN Le et al. (2020) 54.42 39.81 47.11 32.71 45.34 43.88 53.63 39.67 37.08 35.10 42.30 41.56 41.56
HME Fan et al. (2019) 58.28 56.63 57.45 33.71 65.93 54.40 66.12 39.91 40.18 56.89 37.76 48.17 48.17
LAVisH Lin et al. (2023) 52.86 62.72 57.79 38.33 67.47 55.83 78.65 41.48 32.38 62.18 44.05 51.75 51.75
LAViT Yun et al. (2021) 47.01 47.86 47.43 31.39 66.35 48.01 37.21 53.02 36.87 43.05 42.17 42.46 42.46
MCAN Yu et al. (2019) 67.59 54.49 61.04 45.64 64.37 58.62 59.29 53.86 45.02 51.49 46.35 51.20 51.20
MCCD Ma et al. (2024) 75.78 63.43 69.60 61.76 73.43 68.80 76.18 50.55 50.92 62.15 66.95 61.35 61.35
PSAC Li et al. (2019) 54.85 52.77 53.81 37.99 66.83 53.25 53.05 47.14 38.14 48.53 36.46 44.66 44.66
Table 4: Comparison with state-of-the-art methods on the Music-AVQA-R Ma et al. (2024) test set. We report the accuracy for Audio (Counting, Comparative), Visual (Counting, Location), and Audio-Visual (Existential, Counting, Location, Comparative, Temporal) question types, along with the average accuracy for Audio, Visual, Audio-Visual, and overall. Methods highlighted with a gray background incorporate spatial-temporal designs.
@@ Appendix B Representative Examples of Music Performance Scene Types
(a) Methods on Music-AVQA Li et al. (2022).
(b) Methods on Music-AVQA-R Ma et al. (2024).
Figure 5: Radar plots showing the per-type average accuracy of model groups with and without spatial-temporal design across 13 QA categories on (a) Music-AVQA Li et al. (2022) and (b) Music-AVQA-R Ma et al. (2024). Each axis corresponds to a QA type spanning audio, visual, and audio-visual reasoning, including the overall average (Total-Average). The filled  green polygon in Figure 3(a) and  purple polygon in Figure 3(b) represent the mean accuracy across QA types for models with spatial-temporal design, while the blue polygon represents the average performance of models without such design. Models with spatial-temporal design consistently achieve higher accuracy across all modality groups. These advantages persist under distribution shift in the robustness-focused Music-AVQA-R dataset.
@@ Appendix C Representative Examples of Music AVQA Question Types
Appendix D Full-Size Versions of Figures 2 and 3
This Appendix Section presents full-size reproductions of the main-text figures. Specifically, Figure 7 is the enlarged version of Figure 2 (Figure 2, accuracy comparisons by modality), and Figure 5 is the enlarged version of Figure 3 (Figure 3, per-type radar plots). The scale is increased to improve legibility of axes, legends, and model names.
(a) Audio Counting QA.
(b) Audio QA Average.
(c) Visual Counting QA.
(d) Visual Location QA.
(e) Visual QA Average.
(f) Audio-Visual Location QA.
(g) Audio-Visual Comparative QA.
(h) Audio-Visual Temporal QA.
(i) Audio-Visual QA Average.
@@ E.1 Music AVQA Datasets
Table 6: Evolution and characteristics of Music AVQA datasets: a comparative overview of MUSIC-AVQA Li et al. (2022), MUSIC-AVQA v2.0 Liu et al. (2024b), and MUSIC-AVQA-R Ma et al. (2024) benchmarks.
@@
MUSIC-AVQA Li et al. (2022) The MUSIC-AVQA dataset represents a significant contribution to audio-visual question answering research, comprising 9,288 videos with over 150 hours of musical performances covering 22 instruments, generating 45,867 question-answer pairs. The dataset is randomly split into training, validation, and testing sets with 32,087, 4,595, and 9,185 QA pairs respectively, spanning 33 question templates across 9 question types including existential, location, counting, comparative, and temporal questions.
MUSIC-AVQA v2.0 Liu et al. (2024b) The MUSIC-AVQA v2.0 dataset builds upon the original MUSIC-AVQA by addressing data bias issues, comprising 10,518 videos (9,288 from the original plus 1,230 new videos) with musical performances covering 22 instruments, generating approximately 54,000 question-answer pairs. The balanced dataset splits into training, validation, and testing sets with 36,700, 5,250, and 10,819 QA pairs respectively, spanning 33 question templates across 9 question types. The authors specifically balance 15 biased templates by ensuring no dominant answers exceed 60% for binary questions or 50% for multi-class questions, particularly enhancing representation of underrepresented answers in existential, counting, temporal, location, and comparative question categories.
MUSIC-AVQA-R Ma et al. (2024) The MUSIC-AVQA-R dataset proposed in this paper is an extension of MUSIC-AVQA specifically designed to evaluate the robustness of audio-visual question answering models. It expands the original test set through a human-machine collaboration mechanism that rephrases each question 25 times, increasing the number of questions from 9,129 to 211,572, and introduces distribution shifts to categorize questions into head (common) and tail (rare) samples. Compared to the original dataset, MUSIC-AVQA-R features a vocabulary size of 465 (five times larger than MUSIC-AVQA), provides more diverse question formulations while preserving inherent biases in the training and validation sets, and offers three evaluation metrics—head accuracy, tail accuracy, and overall accuracy—enabling researchers to assess model performance in both in-distribution and out-of-distribution scenarios, making it the first dataset specifically designed for robustness evaluation in audio-visual question answering tasks.
@@ E.2 Key Difference from Other AVQA Datasets
Table 7 contrasts the Music-AVQA dataset Li et al. (2022) with several widely-used AVQA benchmarks Yang et al. (2022); Mangalam et al. (2023); Xie et al. (2024a); Liu et al. (2025a); Chen et al. (2020). For each dataset, we highlight the most salient divergence from the music-specific setting, focusing on aspects such as task format, content domain, temporal scope, and the presence or absence of fine-grained musical reasoning.
Table 7: Other representative benchmarks (AVQA Yang et al. (2022), EgoSchema Mangalam et al. (2023), FunQA Xie et al. (2024a), VALOR-1M Liu et al. (2025a), and VGG-Sound Chen et al. (2020)) and the key difference each bears with respect to Music-AVQA Li et al. (2022).
@@
AVQA Yang et al. (2022) Builds multiple-choice QA on everyday VGG-Sound clips; questions target generic activities and causal relations in real-life videos, so it lacks the fine-grained instrument/sound localization and music-theory knowledge required in MUSIC-AVQA.
@@
EgoSchema Mangalam et al. (2023) Uses first-person (Ego4D) footage that is three-minutes long, stressing long-range temporal reasoning in egocentric daily tasks; audio cues are incidental and the task is 5-way multiple choice, very different from the short, professionally filmed music performances and open-ended answers in MUSIC-AVQA.
@@
FunQA Xie et al. (2024a) Focuses on“surprising” humour/creative/magic clips (4.3 k videos, 312 k QAs) that test commonsense violations; audio is often background and questions centre on counter-intuitive visual events, not on synchronised musical notes or instrument semantics.
@@
VALOR-1M Liu et al. (2025a) A pre-training corpus (1 M videos) with tri-modal captions meant for retrieval/captioning; QA supervision is extremely sparse and relies on auto-generated subtitles, so it serves as a foundation model resource rather than a targeted AVQA evaluation set like MUSIC-AVQA.
@@
VGG-Sound Chen et al. (2020) It is an audio-visual correspondent dataset consisting of short clips of audio sounds from YouTube. And it provides raw audio–visual correspondence but no question–answer supervision or fine-grained reasoning labels.
Appendix F How Music AVQA Differs from Traditional Multimodal Tasks
Recent progress in multimodal research has greatly expanded the scope of multimodal modeling, spanning vision–language pretraining Zhang et al. (2025a, 2022); Li et al. (2025); Liu et al. (2025b); Chen et al. (2024a), video captioning Jian et al. (2023), multimodal QA Li et al. (2024b); Diao et al. (2025b), multimodal reasoning Han et al. (2024b); Xie et al. (2026); Diao et al. (2026); Wang et al. (2025a, b, c, 2024b); Zhang et al. (2025b), multimodal content generation and dialogue Gao et al. (2024); He et al. (2025b); Diao et al. (2025a); Jian et al. (2024), retrieval He et al. (2025a); Liu et al. (2026), audio–visual representation learning Gong et al. (2022); Diao et al. (2023), multimodal classification and clustering Qian et al. (2022); Yao et al. (2024b, a); Xie et al. (2024b, 2025), fMRI understanding and decoding Wei et al. (2025b, a). Against this broader backdrop, Music AVQA stands out within audio–visual–text (AVT) tasks for requiring models to disentangle polyphonic, continuous sound streams, bind them to precise visual sources, and reason with explicit musical knowledge. Below, we discuss several representative multimodal task settings as points of comparison.
Generic Audio–Visual Question Answering.
Classic AVQA benchmarks Gemmeke et al. (2017); Chen et al. (2020); Tian et al. (2018); Zhou et al. (2022) present short clips with a single audible event (e.g., a dog bark) and ask coarse “what/where/when” questions that can be answered once the sounding object is localized and its time span identified. The AVQA dataset itself illustrates this design: most videos contain one foreground sound and minimal polyphony, so millisecond-level alignment is unnecessary Li et al. (2022); Yang et al. (2022). Music AVQA, in contrast, poses queries such as “Which violin enters after the flute?”, which requires the model to track multiple overlapping instruments and reason over precise temporal order, a level of granularity generic AVQA never targets.
Video Captioning.
Datasets like MSR-VTT Xu et al. (2016), GIF Li et al. (2016), ActivityNet Captions Krishna et al. (2017), How2 Sanabria et al. (2018), and Vatex Wang et al. (2019) evaluate whether a system can produce one or two fluent sentences that summarize the gist of a clip. Moreover, temporal mis-alignment of a few seconds or the omission of background sounds rarely affects the score. Music AVQA eliminates summarisation entirely and replaces it with beat-accurate, question-driven reasoning: the model must pinpoint note onsets, match them to performers, and compare rhythmic or dynamic patterns tasks far beyond the scope of generic captioning.
Instructional Video QA.
TutorialVQA Colas et al. (2020), YouCookQA Wang et al. (2021) and HowToVQA69M Yang et al. (2021) frame question answering around narrated procedural videos whose audio channel is dominated by speech that explicitly describes each step. The narration acts as a guide-track and overlapping non-speech sounds are rare and not queried. Music AVQA removes narration altogether and treats the dense musical audio as the primary reasoning target, forcing models to infer structure (beats, phrases, instrument entrances) directly from raw sound rather than from textual guidance.
Audio–Visual Scene-Aware Dialogue.
The AVSD dataset Alamri et al. (2019) and its follow-up Alamri et al. (2018); Schwartz et al. (2019); D’Haro et al. (2020); Shah et al. (2022) train systems to hold multi-turn conversations about short ($\approx$ 10 s) household videos, grounding answers in coarse scene context while maintaining dialog coherence. Acoustic events are typically brief (speech, clatter) and alignment at $\pm 1$ s suffices. Music AVQA disregards conversational flow and instead demands hierarchical timing: beats, bars, sections. Every answer depends on tight audio–video synchrony, not on dialog history management. Recent studies on diversified and persona-aware dialogue generation Li and Deng (2023); Li et al. (2024a) further underscore the importance of fine-grained temporal and semantic alignment—even in unimodal dialogue settings.
Multimodal Sentiment Analysis.
Benchmarks such as CMU-MOSI Zadeh et al. (2016) and CMU-MOSEI Tsai et al. (2019) fuse text transcripts, facial expression and prosody to predict sentiment or emotion over 5 to 30 second clips. Overlap between speakers or sound sources is noise to suppress, and no explicit source-binding is required. In Music AVQA overlap is the signal: models must isolate each instrument’s contribution, bind it to its on-screen performer and reason about their relationships (loudness, order, count). Other emotion datasets, like IEMOCAP Busso et al. (2008), MELD Poria et al. (2019), and CH-SIMS Yu et al. (2020), use 5–30 s clips, treat overlapping voices as background noise, and label sentiment at the utterance—rather than source—level.
Cross-Modal Retrieval.
Contrastive systems such as CLIP Radford et al. (2021) (vision–text), ALIGN Jia et al. (2021), LiT Zhai et al. (2022), or ImageBind Girdhar et al. (2023) (audio–vision–text) learn global embeddings and judge success by top-$k$ similarity—minor temporal or spatial errors barely change the score. Music AVQA instead penalises any mis-alignment: swapping left/right instruments or missing a single beat flips the answer. The task therefore demands persistent token-level grounding rather than coarse embedding proximity.
@@ Appendix F How Music AVQA Differs from Traditional Multimodal Tasks
Appendix G Details of Music AVQA Methods with Spatial-Temporal Designs
@@
Amuse Diao et al. (2024) Focuses on music performance scenarios by aligning time segments in both audio and video streams via a cross-attention paradigm. Exploits synchronized features (such as beat-level or note-level alignment) to capture subtle temporal dependencies among instruments in dense music passages. By integrating rhythmic cues and cross-modal interactions, it is particularly suited for questions that involve multiple instruments playing simultaneously or changing their patterns over time.
@@
AVST Li et al. (2022) Proposes a spatio-temporal grounded audio-visual approach that explicitly localizes sounding objects in each frame while applying a question-guided temporal attention mechanism. The model grounds audio-visual events and emphasizes which frames (visual) and which segments (audio) are most relevant for question answering. By combining localized visual features and temporal cues, it captures object interactions over time and can better handle questions involving spatial and temporal reasoning.
@@
CIGN Mo et al. (2023) Learns audio-visual class tokens and an Audio-Visual Continual Grouping module that, at every time-step, pulls together frame-level spectrogram features and region features into category-aware clusters. A token-distillation schedule preserves past knowledge while the regrouping logic tracks objects and sounds through the video’s timeline, giving the model temporally consistent, cross-modal semantics for spatial-temporal reasoning.
@@
DCL Lv et al. (2023) Introduces a Disentangled Counterfactual Learning framework to handle physical audio-visual commonsense reasoning tasks. Decomposes video signals into static (time-invariant) and dynamic (time-varying) factors using a VAE-based encoder, enabling clearer separation of constant background features from changing events. Additionally employs a counterfactual intervention module on the dynamic factors to perform causal reasoning, helping the model answer “what if” questions related to temporal and event relationships.
@@
DG-SCT Duan et al. (2023) Introduces a Dual-Guided Spatial-Channel-Temporal (DG-SCT) attention layer that is injected in every frozen audio and visual transformer block. Audio prompts steer visual tokens (and vice-versa) via bidirectional attention that highlights salient spatial regions, discriminative feature channels, and pivotal temporal segments, producing fine-grained spatio-temporal alignments that boost related tasks.
@@
EEMC Wang et al. (2024a) Divides audio/video into 1-s slices and fuses them with text through a Temporal Bi-modal Transformer backed by a cached-memory mechanism that magnifies sudden changes across time. The resulting multimodal cue stream then serves as a cross-attention prompt for the segmentation decoder, enabling precise localisation of objects and events as their spatial footprints and temporal order evolve.
@@
LAST-Att Liu et al. (2024b) Tackles audio-visual question answering with a repeated cross-attention strategy. Uses Swin-Transformer-v2 for visual frame features and a specialized Audio Spectrogram Transformer for audio, then merges them based on the question. By repeatedly “attending” to the most relevant frames and spectrogram patches, it effectively localizes musical actions (e.g., a pianist’s keystrokes) over time. This design is well suited for intricate temporal queries and locating key audio events in dense musical content.
@@
LAVisH Lin et al. (2023) Adds a lightweight Latent Audio-Visual HYbrid adapter to every layer of a frozen ViT. A compact pool of latent tokens acts as a cross-attention bottleneck, letting audio frames gate visual tokens (and vice-versa) as the video unfolds, so spatial patches and framewise dynamics are fused early while keeping the backbone frozen.
@@
LAViT Yun et al. (2021) Targets 360° videos with a transformer that augments each patch by a quaternion-based spherical coordinate and aligns it with audio via joint contrastive objectives. The spherical embedding plus an auxiliary audio-skewness prediction head lets the model reason about where (on the sphere) and when a sound arises, delivering fine-grained spatial-temporal grounding beyond normal FOV clips.
@@
LSTTA Liu et al. (2023) A parameter-efficient transfer learning approach for audio-visual-language tasks that adds dedicated adapter modules while freezing large pretrained backbones. Splits temporal modeling into two scales: a short-term semantic interaction module (for capturing local correlations such as brief instrumental flourishes) and a long-term semantic filtering module (for broader progressions over many frames). This structure helps the model identify when, how, and for how long different instruments contribute, achieving a refined spatio-temporal representation.
@@
MAVEN Ma et al. (2025) Employs a Multimodal Audio-Visual Epistemic Network that cycles between audio, video and text logits, using debiasing constraints to keep modality-specific and fused predictions consistent over time. The cycle guidance implicitly anchors each question to the correct temporal segments while suppressing spurious correlations.
@@
MCCD Ma et al. (2024) Introduces a Multifaceted Cycle-Collaborative Debiasing objective: KL penalties enlarge the gap between uni-modal and tri-modal logits at every timestep, then force the three unimodal paths to agree with each other. This temporal-cycle training steers attention toward frames (and sounds) that all modalities truly share, yielding stabler spatial-temporal grounding under distribution shift.